the sum of the ages of two brothers is 20. in fiveyears, one is twice the other. find the present ages of the two brothers
Answer by Maths68(1474)

Let
Present age of DISABLED_event_One= a
Present age of Other = b
Sum of the age is 20
a+b = 20 .................(1)
In Fiver years
One will be = a+5 years old
Other will be = b+5 years old
Then
One is twice the other
a+5=2(b+5)
a+5=2b+10
a=2b+10-5
a=2b+5......................(2)
Put the value of a in (1)
2b+5+b = 20
3b+5=20
3b=20-5
3b=15
3b/3=15/3
b=15
Put the value of b in (1)
a+b=20.........(1)
a+15=20
a=20-15
a=5

Present age of DISABLED_event_One= a = 5 years old
Present age of Other = b = 15 years old